OpenStudy (anonymous):

evaluate withput a calculator (2 x 10^3) x (3 x 10^8) (4 x 10^-6) x (2 x 10^5) (3 x 10^5) x (8 x 10^-2)

OpenStudy (anonymous):

6 * 10^11 8*10^-1 = 1/8 = 0.125 24 * 10^3